OpenStudy (anonymous):

solve by elimination pleasee a+b+c=-3 3b-c=4 2a-b-2c=0 how do you work it out

OpenStudy (anonymous):

will take a min

OpenStudy (anonymous):

OK so first we will name the equations to make things simple 1. a + b + c = -3 2. 3b - c = 4 3. 2a-b-2c = 0 First solve for c in equation 2, c = 3b-4 Then plug c into equation 1, a + b +3b - 4 = -3 and solve for a, a = 4b =1 a = 1 - 4b Then plug a and c into equation 3, 2(1-4b) - b -2(3b-4) = 0, solve for b, 2-8b-b-6b+8 = 0 10 - 15b = 0 b = 10/15

OpenStudy (anonymous):

then plug b into the c and a equations, c = 3(10/15) - 4 = 2 -4 = -2 a = 1 -4(10/15) = -25/15

OpenStudy (anonymous):

SO you have a = -25/15, b = 10/15, and c = -2